In 2018, Al, single, had self-employment income of $270,000.

A.) Compute Al's self-employment tax and the amount of his deduction for self-employment taxes.

B.) Compute Al's self-employment tax and the amount of his deduction for self-employment taxes if he had $90,000 of wages and $120,000 of self-employment income.
Answer #1

a. Self employment tax rate is 15.3%, which consists of 12.4% for social security and 2.9% for medicare.

the tax on $270,000 is= 270,000* 0.153= $41,310

b. if the income is $90,000 from wages and $120,000 is from self employment income, then the tax will be:

tax on self employment income= 120,000*0.153= $18,360

plus tax on wages= 90,000*3.07%= $2763

total tax in second case= 41,310+2,763= $44,073